OpenStudy (anonymous):

A wooden pyramid, 9 inches tall, has a square base. A carpenter increases the dimensions of the wooden pyramid by a factor of 4 and makes a larger pyramid with the new dimensions. Describe in complete sentences the ratio of the volumes of the two pyramids

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

the area of a cone is 1/3 * basearea * height

OpenStudy (amistre64):

basearea = side^2 in this case

OpenStudy (amistre64):

factor all the linear measures by 4 1/3 * (4s)^2 * 4*9 = 12*16s^2 the original as: 1/3 * s^2 * 9 = 3*s^2 the ratio of the volume is then 3 to 12*16

OpenStudy (amistre64):

reducing by like factors gets us 1 to 4*16 1 to 64